### v1

# Semantica: An Adaptable Image-Conditioned Diffusion Model

Abstract
We investigate the task of adapting image generative models to different datasets without finetuneing. To this end, we introduce Semantica, an image-conditioned diffusion model capable of generating images based on the semantics of a conditioning image. Semantica is trained exclusively on web-scale image pairs, that is it receives a random image from a webpage as conditional input and models another random image from the same webpage. Our experiments highlight the expressivity of pretrained image encoders and necessity of semantic-based data filtering in achieving high-quality image generation. Once trained, it can adaptively generate new images from a dataset by simply using images from that dataset as input. We study the transfer properties of Semantica on ImageNet, LSUN Churches, LSUN Bedroom and SUN-397.
1 Introduction
ImageNet
Refer to caption      Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
SUN397
Refer to caption Refer to caption
Refer to caption Refer to caption
LSUN Churches
Refer to caption Refer to caption
LSUN Bedroom
Refer to caption Refer to caption
Figure 1: Left: Conditioning image Right: Five samples at guidance 0.5. Semantica is trained exclusively on web-image pairs. During adaptation, it receives a conditioning image and generates samples reflective of semantic information. Semantica requires no label supervision or finetuning.
Machine learning initially focused on optimizing and improving models on small datasets. The field has transitioned to training general purpose models on web-scale data and then finetuning them for specific tasks on smaller datasets. This paradigm shift has lead to state-of-the-art results on a number of different domains. In this paper, we focus on the relatively underexplored task of adapting an image generative model to different datasets. One approach is to simply train a generative model on a large dataset of unlabelled images and finetune them on smaller datasets. While this approach is straight-forward in theory, it requires clever architecture or regularizer design to prevent overfitting in practice (See Sec.2.1). As models scale up, finetuning for every dataset also just becomes increasingly impractical.
Instead, we apply the principle of in-context learning to adapt a image generative model. We train our model called Semantica using image episodes, which are image pairs that belong to the same web page. Therefore, training relies exclusively on the hypothesis that images from the same web page should have some common semantic attributes. For example, it is highly likely that images scraped from a Wikipedia page on dolphins, contain pictures of dolphins. At adaptation, the model receives an image and then generates another image that preserves semantic information.
Semantica consists of a pre-trained image encoder and a diffusion model. The encoder extracts representations from the conditioning image and guides the diffusion model. We thoroughly investigate the impact of different choices for the image encoder (contrastive or self-supervised) and the conditioning mechanisms (global or local). Our experiments also highlight the importance of semantic data filtering to generate high-quality images. Each row in Fig 1 showcases a conditioning image to the left and five generated samples to the right. Semantica is capable of generating high quality and diverse images, reflective of semantic information from the conditioning image across a number of different datasets.
2 Related work
2.1 Generative Transfer
There has been extensive research that studies the adaptation of source-pretrained generative models to a target dataset with adaptation of weights. Initial works study the transfer of discriminators and generators in GANs from a source dataset to a target dataset [47]. Further, [17] show that ImageNet pretraining on a large GAN model is beneficial for transfer to small datasets. A number of papers focus on improving generation quality by adapting only a subset of parameters. These include scale and shift parameters [32], updating only the higher discriminator layers [28], linear combinations of scale and shift parameters [41], modulating kernels or convolutions [59, 58, 10, 2] and singular values [38], mapping networks from noise to latents [46, 29, 53] and latent offsets [12]. Various works apply regularization losses by enforcing constraints to samples/weights by the source generator including elastic weight regularization [27], domain correspondence [33, 16, 22], contrastive learning [60], spatial alignment [51], inversion [49, 23, 44], random masks on discriminators [61] and alignment free spatial correlation [30]. Given the increasing popularity of VQ-VAE and diffusion based models, recent work [43] and [61] explore few-shot finetuning on VQ-VAE tokens and diffusion models. We defer to the survey [1] for a detailed exposition of all these methods. In contrast to these works, we explore training a generator on web-scale images and study their transfer to standard small-scale image datasets. Retrieval augmented models [7, 5] compute nearest neighbours for a query image across a bank of memory images. These retrieved neighbors facilitate the training or generation process. Unlike these methods, we do not require access to a memory bank of images during train or test time.
2.2 Diffusion
Diffusion and score-based generative models have become increasingly successful in modelling images [20, 40, 31, 4], videos [18, 42] and audio [25]. As generation quality has steadily improved, they have been used in contexts with more and more conditioning variables. Well-known examples are text-to-image and text-to-video modelling, where the conditioning variable is text. In this case, the conditioning variable can be seen as a sequence from which cross-attention layers communicate to the feature maps of the image or video, i.e. what the diffusion model is learning to generate [40, 31]. As the desire for controllable generation increases, solutions such as ControlNet [56] have been developed. ControlNet takes in conditioning images of the same size as the generations, and uses a copy of the UNet to learn an encoder for the conditioning signals. Although this encoder trains fast due to parameter initialization from a pretrained diffusion UNet, it is difficult to deal with different sized inputs. In those cases, only conditioning via cross-attention as done in [52] overcomes the in-place additions between the ControlNet encoder and the base UNet. Conditioning on images as context has produced impressive results, turning scribbles or edge detections into high quality image generations [48] and discriminative tasks [3, 26]. In contrast with the above mentioned techniques, our framework relies on general web-based pretraining for semantic-based adaptive image generation. Finally, while [13] study the transfer of few-shot diffusion models between small datasets (CIFAR-100 -> miniImageNet, we see in Sec. 5.6 that this can lead to sub-optimal transfer.
3 Background
3.1 Diffusion
Diffusion models learn to generate examples by gradually denoising a diffusion process. For a single datapoint, their loss can be expressed as a squared error between the original datapoint and its prediction:
$\mathbb{E}_{t\sim\mathcal{U}(0,1),{\bm{{\epsilon}}}\sim\mathcal{N}(0,{\mathbf{I}})}\big{[}w(t)||{\bm{x}}-f({\bm{z}}_{t})||^{2}\big{]}\text{ where }{\bm{z}}_{t}=\alpha_{t}{\bm{x}}+\sigma_{t}{\bm{{\epsilon}}}_{t}$ (1)
It is helpful to define $\mathrm{SNR}(t)=\alpha_{t}^{2}/\sigma_{t}^{2}$. In the case of $w(t)=\mathrm{SNR}(t)$ the loss above is equivalent to a loss in ${\bm{{\epsilon}}}$-space, the simple loss from [19]. After training, the denoising model generates samples by taking small steps. Starting at $t=1$ with initial Gaussian noise and one slowly denoises for timesteps $t=1,1-1/N,\ldots$ where the number of sampling steps is $N$. Although many samplers are possible, in this paper we use the standard DDPM sampler [19].
Architecture
In early days, diffusion literature typically used UNets that consisted of ResNet blocks, with optional self-attention layers. More recent architecture either use full Transformers (DiT [35], StableDiffusion) or transformer backbones UViTs [21]. This transformer backbone makes it especially easy to use conditioning signals in these architecture via cross-attention layers. To be precise, one can say that the denoising neural network takes in ${\bm{z}}_{t}\in\mathbb{R}^{H\times W\times 3},t\in\mathbb{R},y\in\mathbb{R}^{T_{y}\times D_{y}}$. In principle it does not matter which diffusion model we use to generate images. In fact, every generative model would suffice. In practice we use the simple diffusion framework [21] because it can learn to generate high resolution images end-to-end without the need of a separate autoencoder.
4 Method
4.1 Episodic Dataset
The training dataset for our model is a subset of WebLI [9] that consists of roughly one billion images from the web. In particular, we use Episodic WebLI [8], where each episode contains randomly sampled loosely related images (i.e., they are clustered according to their URL). We randomly sample an image as the conditioning input and another image from the same episode as the ground-truth "target" image. We do not use any text inputs and only rely on image conditioning.
4.2 Image Encoder
Training an image-conditioned diffusion model requires two components: 1) An image encoder that extracts semantic information from a conditioning image. 2) A diffusion model that denoises a noisy target image conditioned on the image representations. We could train a separate Vision Transformer or ResNet model end-to-end as an image encoder with the diffusion model to learn useful conditioning representations.
Instead, we leverage pre-trained image encoders and condition our diffusion model on their “frozen" representations. This offers two advantages. First, we can precompute representations for all images in the dataset that eliminates expensive forward passes through the encoder during training. Second, by relying on pretrained encoders, we can just focus on scaling the diffusion model.
We investigate ViT image encoders trained with two pretraining strategies: training on web-text image pairs [37] (we use the SigLIP [55]), and self-supervision with teacher pseudo-labels (DINO) [6, 34]. Furthermore, we explore two conditioning strategies to incorporate conditioning information:
  1. 1.
    Token level conditioning: The diffusion model cross-attends to a sequence of tokens from the final transformer block of the image encoder.
  2. 2.
    Global feature conditioning: The encoder provides a single global feature representation from the encoder, being the CLS token for DINO and the output of the MAP block for SigLIP. We normalize and embed this with a dense projection and add it to the noise embedding. This subsequently conditions the diffusion model via FiLM layers [36] as done in standard label-conditioned diffusion.
4.3 Data Filtering
Each episode consists of images that are loosely related, whereas our model assumes conditioning and target images share semantic information. This mismatch may lead the model to waste capacity on modeling irrelevant noisy conditioning-target pairs. To address this, we filter out pairs with low semantic similarity. The pretrained encoder (DINO/SigLIP) computes the global CLS representation from the conditioning and target image. We compute the cosine similarity between the global conditioning/target representations and filter out pairs below a lower threshold. Additionally, to ensure generation of interesting images, we also filter conditioning-target pairs with similarity above a high threshold. This may ensure that generated images retain the semantics of the conditioning image, while being sufficiently different.
5 Experiments
5.1 Evaluation Datasets
We use $10\%$ of the Episodic WebLI dataset for the architecture design ablations and we use the entire dataset for the final model. We evaluate the model by conditioning it on ImageNet validation set images and then computing the FID between the generated samples and the ImageNet validation set. We investigate the transfer of Semantica on four datasets: ImageNet [39], LSUN-Outdoor Church, LSUN-Bedroom [54] and SUN-397 [50]. We report the FID for each dataset using 50K conditioning images randomly sampled from within the dataset. We use the validation dataset for ImageNet while the train set for all other datasets. To assess diversity, we compute the mean LPIPS [57] score between 50K generated samples and their corresponding conditioning images.
5.2 Image Encoder
Here, we evaluate the different encoder and conditioning design choices (Sec. 4.2). We train baseline diffusion models with two frozen image encoders (DINO-v2 B/14 and SigLIP B/16) and two conditioning strategies (cross-attention and FILM).
Refer to captionRefer to captionRefer to caption
Figure 2: We plot ImageNet FID as a function of number of training steps on Episodic WebLI. Left: FID of two image encoders DINO-v2 B/14 and SigLIP B/16 using two different conditioning methods: 1) All encoder output tokens via cross-attention (CA) 2) Global feature representation via FILM. Center: FID of DINO-v2 B/14 + Cross Attention with and without data filtering. Right: FID of DINO-v2 B/14 + Cross Attention with 1x1 patching and higher learning rate.
Fig. 2 left shows ImageNet FID as a function of number of training steps on Episodic WebLI. Token-level conditioning significantly outperforms global feature conditioning for both DINO-v2 and SigLIP, highlighting the importance of low-level features for this particular task. While SigLIP FiLM performs better than DINO-v2 FiLM, this trend reverses with cross-attention. This suggests a greater compatibility of DINO-v2 low-level features over CLIP low-level features for semantics-based image generation.
For this reason, we will now use DINO-v2 frozen representations with a cross-attention conditioning mechanism as our main architecture.
5.3 Data Filtering
We investigate the impact of semantic data filtering, as described in Sec. 4.3. We heuristically set the lower and higher thresholds to 0.3 and 0.9 respectively. Fig. 2 middle shows the FID of the DINO-v2 B/14 + Cross Attention with and without data filtering. Similarity-based data filtering in DINO feature space positively impacts the generation quality and improves FID by greater than 10. In future work, we can explore tuning these thresholds for a desired quality-diversity tradeoff or even directly conditioning the diffusion model on the desired similarity with the conditioning image.
5.4 Hyperparameters
For our baseline diffusion model, we inherit all hyper-parameters from the ImageNet label conditioned diffusion model. The denoising model follows a U-ViT architecture that operates on $256\times 256$ images. The architecture consists of a initial $2\times 2$ patchification layer [11, 45] that downsamples the noised image by a factor of 2. The model has four downsampling stages, where each stage downsamples the feature maps by a factor of 2 at its output. The resolution of the lowest feature map is $16\times 16$. Transformer blocks operate at the stages with the two lowest resolutions $16\times 16$ and $32\times 32$ and convolutional blocks operate in the remaining stages. The four downsampling stages have 128, 256, 512 and 1024 channels. The first three stages have three blocks each and the last stage has sixteen blocks. The optimizer is Adam [24] with parameters $\beta_{1}=0.9$, $\beta_{2}=0.99$, $\epsilon=1e^{-12}$, batch size of 2048 and a learning rate of $1e^{-4}$. We also use Polyak averaging with a decay factor of $0.9999$. The diffusion loss parameters include v-prediction with loss in epsilon phase and a cosine adjusted schedule with a noise resolution of 32. We use the DDPM sampler with an interpolation of 0.2 (standard deviation is $\sigma_{t\to s}^{0.2}\sigma_{st}^{0.8}$) and no guidance for our ablations. Each ablation run utilizes 256 TPUv3 [15] chips around 2 days. However, the consistent ranking of different ablations throughout training can allow for a much shorter training schedule to identify the best model.
We finally perform a lightweight sweep of the diffusion model hyper-parameters. Notably, two minor modifications improve generation quality: 1) Replacing the 2x2 patching with no patching. The model consists of a $1\times 1$ convolution at the input and output. The downsampling proceeds with an additional stage consisting of two convolutional blocks and 128 channels each. 2) Increasing the learning rate from $41e-$ 4 to $42e-$ 4. Fig. 2 right shows the improvement in FID of the DINO B/14 encoder with these modifications. Other changes such as tuning noise schedules, dropout rates or Polyak decay factors cause negligible changes or decrease sample quality.
5.5 Final Training
In our final training phase, we make the following four adjustments:
  1. 1.
    Replace DINO-v2 ViT-B/14 conditioning features with the larger DINO ViT-L/14 conditioning features.
  2. 2.
    Scale the model by doubling the number of channels across all UViT stages. The only exception is the compute-heavy first downsampling stage which we keep at 128.
  3. 3.
    Expand training data to include the entirety of episodic WebLI which was previously $10\%$ in the ablations.
  4. 4.
    Extend the training duration from 300K steps to two million steps
We train our model on 256 Cloud TPUv5 chips [14] for approximately 12 days with a constant learning rate of $2\text{e}{-4}$.
5.6 Label Grouped Baseline
Here we compare Semantica to a baseline that has direct label supervision (for example, ImageNet). Recall that the conditioning and target image belong to the same webpage. However, in the presence of label supervision (as in ImageNet), the target and conditioning image can just belong to the same class label. So as a supervised baseline, we group images on ImageNet as per their label and train Semantica on this dataset. Table 1 compares the FID of the Label Grouped baseline (LG) to Semantica. Since LG is trained on ImageNet, it significantly outperforms Semantica (FID 4.8 vs FID 18.4). However, this trend reverses on all other datasets, where Semantica outperforms LG. Both the supervised baseline and Semantica rely on the DINOv2 encoder which was trained on a wide variety of data sources. Therefore the encoder itself may provide useful representations on a number of datasets. But training LG just on ImageNet, might limit the diffusion model’s exposure from non ImageNet images, potentially explaining its significant performance drop on all other datasets.
5.7 One Shot Results
We evaluate Semantica’s one-shot image generation capabilities. Each row in Fig. 1 showcases a conditioning image to the left and five generated samples to the right. We observe that Semantica is capable of generating high quality and diverse images.
Table 1 reports quantitative results. Semantica receives 1000 and 397 conditioning images for ImageNet and SUN397 respectively, one per class. The model generates 50 and 120 samples per class for ImageNet and SUN397 totalling 50000 and 47840 generated samples. We then report the FID with 50000 ground truth samples. Semantica achieves a FID of 19.1 and 9.5 on ImageNet and SUN397 respectively. On increasing the guidance to 0.5, the FID improves to 9.3 and 7.1 respectively. Simply replicating the conditioning image Copy 50 and 120 times, achieves a much worse FID of 38.4 and 70.6.
5.8 Effect of Guidance
We investigate the quality/diversity tradeoff offered by Semantica’s guidance parameter. Fig. 3 Left illustrates the FID and LPIPS on ImageNet for different guidance factors. Note that we filter out near duplicates with a semantic similarity greater than 0.9 from the training set. This makes it unlikely that the model copies the conditioning image exactly to achieve a perfect FID. With no guidance, Semantica achieves a FID of 18.4 on ImageNet. The high LPIPS score 0.72 further reflects the model’s ability to generate diverse samples. On increasing the guidance upto a certain extent, we tradeoff quality with diversity. Notably, at guidance factor 0.5, which is commonly used in label conditioned generation, Semantica achieves a FID of 6.2 with a LPIPS score of 0.68. Fig. 4, displays five conditioning images from ImageNet and the generated samples at different guidance factors. At guidance factor 0.0, the samples reflect a broad semantic category from the conditioning image. Increasing the guidance factor leads to samples that incorporate more specific details from the conditioning image. For example, with the conditioning image of the dog and the kid, Semantica stars with a sample of a dog. The specific breed of the dog and the child in the image appear as we amplify the guidance.
Refer to captionRefer to caption
Figure 3: LPIPS (Higher is more diverse) vs FID (Lower is better) for five different datasets (Left: ImageNet, Right: LSUN Bedroom, LSUN Churches, SUN397). Each point represents a different guidance factor. As expected, we tradeoff quality with diversity on increasing guidance, but even with lower diversity the model does not collapse completely to its conditioning image.
ImageNet Bedroom Church SUN397
Label grouped 4.8 46.2 27.1 29.7
Semantica 18.4 6.2 17.3 6.7
Guidance @@@@ 0.5
Label grouped 5.1 34.2 20.4 22.4
Semantica 6.2 2.4 4.0 2.5
ImageNet SUN397
Semantica 19.1 9.5
+ Guidance @@@@ 0.5 9.3 7.1
Copy 38.4 70.6
Table 1: Left: Comparison between Semantica and a Label Grouped baseline (LG) trained on ImageNet. The conditioning and target pairs have the same label for LG. LG outperforms Semantica in-distribution and performs worse on out-of-distribution datasets. Right (One Shot Generation Results): Evaluates Semantica one-shot image generation capabilities. Semantica receives 1000 and 397 conditioning images for ImageNet and SUN397 respectively, one-per-class. The model generates 50 and 120 samples per-class for ImageNet and SUN397 totalling 50000 and 47840 generated samples. We then report the FID with 50000 ground truth samples. Copy is a baseline where the same per-class conditioning image is just replicated 50 and 120 times respectively.
Figure 4: Left: Conditioning Image from ImageNet. Right: Generated samples with guidance factors 0.0, 0.1, 0.2, 0.5 and 1.0. At guidance factor 0.0, the samples reflect a broad semantic category from the conditioning image. Increasing the guidance factor leads to samples that incorporate more specific details from the conditioning image.
Fig. 3 Right demonstrates quantitative results on three smaller datasets LSUN Bedroom, LSUN Church and SUN-397. We observe a similar tradeoff between quality and diversity as we vary the guidance factor. With zero guidance, Semantica achieves a FID of 6.2, 17.3 and 6.7 on LSUN Bedroom, LSUN Church and SUN397 respectively. On increasing the guidance to 0.5, the FID improves to 2.4, 4.0 and 2.5 respectively.
Fig. 5 showcases samples for each small dataset across various guidance factors. In row four, the bed (main object) persists across all guidance levels, while the chair and the fence appear at high guidance levels. Row five exhibits a similar effect: the number of minarets in the generated church increases from one to two and the shape of the main dome begins to resemble the conditioning image. In row one, the sample resembles toys with zero guidance, the sample resembles toys but transforms into a crowded convention as guidance increases.
SUN397
Refer to caption Refer to caption
Refer to caption Refer to caption
LSUN Bedrooms
Refer to caption Refer to caption
Refer to caption Refer to caption
LSUN Churches
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 5: Left: Conditioning Images from SUN397 (Top two rows), LSUN Bedrooms (Middle two rows) and LSUN churches (Last two rows). Right: Generated samples with guidance factors 0.0, 0.1, 0.2, 0.5, 1.0 and 1.5.
6 Conclusion
We present Semantica, an image-conditioned diffusion model capable of generating high-quality and diverse images based on the semantics of a conditioning image. Our pretraining strategy is simple, the diffusion model models a target image with the conditioning image coming from the same webpage. Our experiments also further emphasize the expressivity of pretrained image encoders and necessity of semantic-based data filtering.
Limitations
We identify the following limitations of Semantica. 1) While adapting Semantica to downstream datasets, requires no finetuneing, a limitation is that training Semantica requires high computational resources as is the case with most generative models. However, any improvements that make diffusion models more efficient is likely to transfer to Semantica. 2) Our model depends on a frozen encoder and we do not modify the weights. With more resources, we can also study the tradeoffs with this design decision. 3) With a very high guidance factor (>1.0) , Semantica might converge to an oversaturated conditioning image. An area of improvement would be to find ways to steer it to another image with the same semantic information. 4) Our model as is does not incorporate additional conditioning signals. Generalizing the model to utilize additional conditioning information can be an exciting area of future research. 5) While our model generally produces diverse and high-quality images, sometimes artifacts can appear especially at guidance 0.0 and in areas like lower-level structure and human faces. Scaling our model further may remove these artifacts.
Appendix A Broader Impact Statement
This paper introduces a model capable of generating images that reflect the semantic information of a given input image. Potential positive applications include being a creative aid, augmentation of an image retrieval system or enhancing the size of an image dataset. While the model lacks direct controllability thus mitigating some negative applications like creation of personalized deepfakes, our model can still generate deepfakes. Therefore, we advocate for responsible usage for our model.
Appendix B Standard Errors
In general, the variance in FID is small. We compute the variance across different random seeds and different randomly sampled one shot datasets on ImageNet. With a fixed one-shot dataset and five different random seeds, we report a FID of $9.4\pm 0.02$. We also sample five different one-shot datasets and report a FID of $9.5\pm 0.06$.
Appendix C ImageNet Samples
Figure 6: Set 1: Conditioning Images from ImageNet
Figure 7: Set 1: Random Samples 1
Figure 8: Set 1: Random Samples 2
Figure 9: Set 1: Random Samples 3
Figure 10: Set 2: Conditioning Images from ImageNet
Figure 11: Set 2: Random Samples 1
Figure 12: Set 2: Random Samples 2
Figure 13: Set 2: Random Samples 3
Figure 14: Set 3: Conditioning Images from ImageNet
Figure 15: Set 3: Random Samples 1
Figure 16: Set 3: Random Samples 2
Figure 17: Set 3: Random Samples 3
Appendix D LSUN Bedroom Samples
Figure 18: Conditioning Images from LSUN Bedroom
Figure 19: Random Samples 1
Figure 20: Random Samples 2
Figure 21: Random Samples 3
Appendix E SUN 397 Samples
Figure 22: Conditioning Images from SUN397
Figure 23: Random Samples 1
Figure 24: Random Samples 2
Figure 25: Random Samples 3
Appendix F LSUN Church Samples
Figure 26: Set 1: Conditioning Images from LSUN Churches
Figure 27: Random Samples 1
Figure 28: Random Samples 2
Figure 29: Random Samples 3

### v3

@@
-Semantica: An Adaptable Image-Conditioned Diffusion Model
+Conditional Diffusion on Web-Scale Image Pairs leads to Diverse Image Variations
 Abstract
-We investigate the task of adapting image generative models to different datasets without finetuneing.
-To this end, we introduce Semantica, an image-conditioned diffusion model capable of generating images based on the semantics of a conditioning image.
-Semantica is trained exclusively on web-scale image pairs, that is it receives a random image from a webpage as conditional input and models another random image from the same webpage.
-Our experiments highlight the expressivity of pretrained image encoders and necessity of semantic-based data filtering in achieving high-quality image generation.
-Once trained, it can adaptively generate new images from a dataset by simply using images from that dataset as input.
-We study the transfer properties of Semantica on ImageNet, LSUN Churches, LSUN Bedroom and SUN-397.
+Generating image variations, where a model produces variations of an input image while preserving the semantic context has gained increasing attention.
+Current image variation techniques involve adapting a text-to-image model to reconstruct an input image conditioned on the same image.
+We first demonstrate that a diffusion model trained to reconstruct an input image from frozen embeddings, can reconstruct the image with minor variations.
+Second, inspired by how text-to-image models learn from web-scale text-image pairs, we explore a new pretraining strategy to generate image variations using a large collection of image pairs.
+Our diffusion model Semantica receives a random (encoded) image from a webpage as conditional input and denoises another noisy random image from the same webpage.
+We carefully examine various design choices for the image encoder, given its crucial role in extracting relevant context from the input image.
+Once trained, Semantica can adaptively generate new images from a dataset by simply using images from that dataset as input.
+Finally, we identify limitations in standard image consistency metrics for evaluating image variations and propose alternative metrics based on few-shot generation.
 1 Introduction
+Machine learning initially focused on optimizing and improving models on small datasets.
+The field has transitioned to training general purpose models on web-scale data and then finetuning them for specific tasks on smaller datasets.
+This paradigm shift has lead to state-of-the-art results on a number of different domains.
+In this paper, we focus on the relatively underexplored task of adapting an image generative model to different datasets.
+One approach is to simply train a generative model on a large dataset of unlabelled images and finetune them on smaller datasets.
+While this approach is straight-forward in theory, it requires clever architecture or regularizer design to prevent overfitting in practice (See Sec.2.2).
+As models scale up, finetuning for every dataset also just becomes increasingly impractical.
+Semantica
+IP-Adapter
+SD v2 - IV
+Semantica
+IP-Adapter
+SD v2 - IV
+Figure 1: Each grid presents a conditioning image at the top followed by $512\times 512$ image variations generated by Semantica, IP-Adapter, and SDv2 IV.
+Samples generated by a semantic image-variation model should maintain semantic consistency with the conditioning image while also being sufficiently diverse.
+Semantica demonstrates greater diversity than IP-Adapter while preserving semantic context.
+While SD v2 generates diverse outputs, the generated outputs are often not congruent with the context image.
+Additional samples are present in App.
+A.
+Image-conditioned diffusion models are now increasingly used to adapt generative models to new datasets, also known as image variations (Ye et al., 2023; Pinkney, 2022; Xu et al., 2023b).
+First, an image encoder trained on a different upstream task, such as self-supervised learning (DINO (Oquab et al., 2023)) or contrastive-learning on web-image text pairs (CLIP (Radford et al., 2021)) produces frozen embeddings.
+The frozen embeddings then condition a diffusion model usually pretrained on text-to-image, which is finetuned to reconstruct the original image.
+However in these models, the study of image variations often remains a secondary objective.
+In this paper, we take a step back and directly analyze image variations in isolation.
+To avoid ambiguity involved in generating multiple objects in an image, we focus our evaluation on datasets with a single dominant object.
+We start with a vanilla image-conditional diffusion architecture that is composed of a frozen image encoder and conditions the diffusion model with cross-attention.
+As done in prior works, we train the diffusion model to reconstruct images from frozen embeddings.
+We demonstrate that without text-to-image pretraining or co-training, this model qualitatively achieves near-perfect image reconstruction.
+This suggests that the capacity to generate image variations via reconstruction is due to the implicit regularization inherent in the pre-trained or co-trained text-to-image model.
+While empirically this may be sufficient to generate plausible image variations, the relationship between the text-to-image model, image variations and scale remains unclear.
+In this paper, we explore a different pretraining strategy with the same image-conditioned diffusion architecture.
+We train our model Semantica using image episodes, which are image pairs that belong to the same web page.
+Therefore, training relies exclusively on the hypothesis that images from the same web page should have some common semantic attributes.
+For example, it is probable that images scraped from a Wikipedia page on dolphins, contain pictures of dolphins.
+To generate image variations, the model receives an image and then generates another image that preserves semantic information.
+Under this pre-training strategy, our experiments demonstrate that scaling both the image encoder and the diffusion decoder steadily improve image variation quality.
+In Fig.
+1, we compare Semantica to state-of-the-art image variation models.
+Semantica is capable of generating high quality and diverse images, reflective of semantic information from the conditioning image.
+Evaluating image variations is non-trivial.
+Unlike standard image generative modeling where a model generates images from scratch, a model has access to the entire test set of images via conditioning when generating image variations.
+This means a model could simply copy the conditioning image and achieve high scores both at distribution-level and instance-level metrics.
+To bridge this limitation in existing metrics, we instead propose to evaluate image-variations exclusively in the few-shot setting.
+Concretely, we limit the number of conditioning images available to the model and then measure its ability to model the test distribution.
+Our main contributions are:
+  • Current techniques train diffusion models on image reconstruction to produce image variations.
+Our analysis shows that diffusion models trained to reconstruct images from frozen embeddings produce only minor low-level variations.
+  • We explore an alternative pretraining strategy for generating image variations.
+This involves conditioning a diffusion model with a random image from a webpage and training it to denoise a different random image from the same webpage.
+  • We rigorously compare DINOv2 and CLIP as frozen image encoders to produce image variations.
+While CLIP is now the standard image encoder, our experiments demonstrates that DINOv2 yields superior performance.
+  • Standard image-level metrics such as LPIPS and distribution-level metrics such as FID fail to capture diversity in image variations.
+To address this, we introduce few-shot metrics designed to assess the diversity in image variations.
+2 Related Work
+2.1 Image Variations
+SD-V2 Image Variations (Pinkney, 2022), IP-Adapter (Ye et al., 2023), MultiFusion (Bellagente et al., 2024) and Verstatile Diffusion (Xu et al., 2023b) generate image variations via image reconstruction.
+Specifically, SD-V2, IP-Adapter and MultiFusion adapt a pretrained Stable Diffusion model.
+SD-V2 swaps the frozen CLIP text embedding with the CLIP image embedding, first only finetunes the cross-attention layers of the SD model to attend to the image embedding and then the entire backbone.
+IP-Adapter trains a adapter layer to the output of the clip image embedding and additional decoupled cross-attention layers.
+MultiFusion finetunes a LLM to accept additional image inputs.
+The resultant LLM embeddings than condition a pretrained Stable-Diffusion model.
+Versatile Diffusion trains a single model to perform both text-to-image and image variations, with some decoupled components such as cross-attention.
+All these methods use the same input image as the target image, and rely on regularization for variations.
+Bordes et al.
+(2021) analyze the reconstructions generated by diffusion models conditioned on just the global embedding from self-supervised models and show that they can reconstruct image semantics.
+Here we show, with cross-attention based conditioning, near perfect reconstruction can be achieved.
+2.2 Generative Transfer
+Prior to image variations, there has been research that studies the adaptation of source-pretrained generative models to a target dataset with adaptation of weights.
+Initial works study the transfer of discriminators and generators in GANs from a source dataset to a target dataset (Wang et al., 2018).
+Further, Grigoryev et al.
+(2022) show that ImageNet pretraining on a large GAN model is beneficial for transfer to small datasets.
+A number of papers focus on improving generation quality by adapting only a subset of parameters.
+These include scale and shift parameters (Noguchi & Harada, 2019), updating only the higher discriminator layers (Mo et al., 2020), linear combinations of scale and shift parameters (Shahbazi et al., 2021), modulating kernels or convolutions (Zhao et al., 2022a; 2020; Cong et al., 2020; Alanov et al., 2022) and singular values (Robb et al., 2020), mapping networks from noise to latents (Wang et al., 2020; Mondal et al., 2023; Yang et al., 2023) and latent offsets (Duan et al., 2024).
+Various works apply regularization losses by enforcing constraints to samples/weights by the source generator including elastic weight regularization (Li et al., 2020), domain correspondence (Ojha et al., 2021; Gou et al., 2023; Hou et al., 2022), contrastive learning (Zhao et al., 2022b), spatial alignment (Xiao et al., 2022), inversion (Wu et al., 2022; Kato et al., 2023; Thopalli et al., 2023), random masks on discriminators (Zhu et al., 2022) and alignment free spatial correlation (Moon et al., 2023).
+Given the increasing popularity of VQ-VAE and diffusion based models, recent work (Sohn et al., 2023) and (Zhu et al., 2022) explore few-shot finetuning on VQ-VAE tokens and diffusion models.
+We defer to Abdollahzadeh et al.
+(2023) for a detailed exposition of all these methods.
+In contrast to these works, we explore training a generator on web-scale images and study their transfer to standard small-scale image datasets.
+Retrieval augmented models (Casanova et al., 2021; Blattmann et al., 2022) compute nearest neighbours for a query image across a bank of memory images.
+These retrieved neighbors facilitate the training or generation process.
+Unlike these methods, we do not require access to a memory bank of images during train or test time.
+2.3 Diffusion
+Diffusion and score-based generative models have become increasingly successful in modelling images (Ho et al., 2022b; Saharia et al., 2022; Nichol et al., 2022; Balaji et al., 2022), videos (Ho et al., 2022a; Singer et al., 2022) and audio (Kong et al., 2020).
+As generation quality has steadily improved, they have been used in contexts with more and more conditioning variables.
+Well-known examples are text-to-image and text-to-video modelling, where the conditioning variable is text.
+In this case, the conditioning variable can be seen as a sequence from which cross-attention layers communicate to the feature maps of the image or video, i.e.
+what the diffusion model is learning to generate (Saharia et al., 2022; Nichol et al., 2022).
+As the desire for controllable generation increases, solutions such as ControlNet (Zhang et al., 2023) have been developed.
+ControlNet takes in conditioning images of the same size as the generations, and uses a copy of the UNet to learn an encoder for the conditioning signals.
+Although this encoder trains fast due to parameter initialization from a pretrained diffusion UNet, it is difficult to deal with different sized inputs.
+In those cases, only conditioning via cross-attention as done in (Xu et al., 2023a) overcomes the in-place additions between the ControlNet encoder and the base UNet.
+Conditioning on images as context has produced impressive results, turning scribbles or edge detections into high quality image generations (Wang et al., 2023; Najdenkoska et al., 2023) and discriminative tasks (Bai et al., 2023; Li et al., 2023).
+In contrast with the above mentioned techniques, our framework relies on general web-based pretraining for semantic-based adaptive image generation.
+While (Giannone et al., 2022) study the transfer of few-shot diffusion models between small datasets (CIFAR-100 $\rightarrow$ miniImageNet, we see in Sec.
+2 that this can lead to sub-optimal transfer.
+(Liu et al., 2023) employ test-time guidance using similarity scores with a reference image, to steer unconditional generative models.
+However, they still require training a separate unconditional generative model for each domain.
+3 Model
+3.1 Diffusion
+Diffusion models learn to generate examples by gradually denoising a diffusion process.
+For a single datapoint, their loss can be expressed as a squared error between the original datapoint and its prediction:
+$\mathbb{E}_{t\sim\mathcal{U}(0,1),{\bm{{\epsilon}}}\sim\mathcal{N}(0,{\mathbf{I}})}\big{[}w(t)||{\bm{x}}-f({\bm{z}}_{t},t,t)||^{2}\big{]}\text{ where }{\bm{z}}_{t}=\alpha_{t}{\bm{x}}+\sigma_{t}{\bm{{\epsilon}}}_{t}$ (1)
+It is helpful to define $\mathrm{SNR}(t)=\alpha_{t}^{2}/\sigma_{t}^{2}$.
+In the case of $w(t)=\mathrm{SNR}(t)$ the loss above is equivalent to a loss in ${\bm{{\epsilon}}}$-space, the simple loss from Ho et al.
+(2020).
+After training, the denoising model generates samples by taking small steps.
+Starting at $t=1$ with initial Gaussian noise and one slowly denoises for timesteps $t=1,1-1/N,\ldots$ where the number of sampling steps is $N$.
+Although many samplers are possible, in this paper we use the standard DDPM sampler (Ho et al., 2020).
+3.2 Image Encoder
+Training an image-conditioned diffusion model requires an image encoder that extracts semantic information from a conditioning image.
+We could train a separate ViT end-to-end as an image encoder with the diffusion model to learn useful conditioning representations.
+Instead, we leverage pre-trained image encoders and condition our diffusion model on their “frozen” representations.
+This offers two advantages.
+First, we can precompute representations for all images in the dataset that eliminates expensive forward passes through the encoder during training.
+Second, we can use different scales of readily available pre-trained encoders and just focus on scaling the diffusion model.
+We investigate ViT image encoders trained with two pretraining strategies, contrastive learning (SigLIP) (Zhai et al., 2023) and self-supervised learning (DINOv2) (Caron et al., 2021; Oquab et al., 2023).
+3.3 Diffusion Decoder
+In early days, diffusion literature typically used UNets that consisted of ResNet blocks, with optional self-attention layers.
+More recent architecture either use full Transformers (DiT (Peebles & Xie, 2023), StableDiffusion) or UNeTs with transformer backbones (UViTs) ((Hoogeboom et al., 2023).
+The transformer backbone makes it especially easy to use conditioning signals in these architecture via cross-attention layers.
+To be precise, the denoising neural network takes in a noised image at a certain timestep ${\bm{z}}_{t}\in\mathbb{R}^{H\times W\times 3}$, timestep $t\in\mathbb{R}$ and contextual information $c\in\mathbb{R}^{T_{c}\times D_{c}}$.
+In principle it does not matter which diffusion or generative model we use to generate images.
+In practice we use the simple diffusion framework (Hoogeboom et al., 2023) because it can learn to generate high resolution images end-to-end without the need of a separate autoencoder.
+3.4 Conditioning Information
+Formally, the image encoder encodes the context image ${\bm{X}}_{c}\in\mathbb{R}^{H\times W\times C}$ into a sequence of tokens ${\bm{X}}_{C}\in\mathbb{R}^{T_{c}\times D_{c}}$, We follow the encoder-decoder framework in the original Transformer (Vaswani, 2017) to condition the diffusion decoder with context tokens.
+We employ conditioning only in the lowest-resolution transformer in UViT.
+Every self-attention block in the transformer backbone is followed by a cross-attention block, where the diffusion decoder cross-attends to ${\bm{X}}_{c}$.
+In addition to conditioning via cross-attention, we also explore conditioning only using global features using the CLS token.
+Specifically, we normalize the CLS token, embed it with a dense projection and add it to the timescale embedding.
+The resultant embedding then conditions the diffusion model via FiLM layers (Perez et al., 2018).
+4 Image Variations via Reconstruction
+A common technique to generate image variations is to incorporate image-specific context into the denoising objective using frozen embeddings.
+$\mathbb{E}_{t\sim\mathcal{U}(0,1),{\bm{{\epsilon}}}\sim\mathcal{N}(0,{\mathbf{I}})}\big{[}||{\bm{x}}-f({\bm{z}}_{t},t,{\bm{x}}_{c})||^{2}\big{]}$ (2)
+where $f$ is a pre-trained model on a different objective, for example text-to-image modeling (Ye et al., 2023; Bellagente et al., 2024), ${\bm{x}}$ is an image and ${\bm{x}}_{c}$ are frozen embeddings from the same image.
+The objective amounts to reconstruction where the diffusion model is trained to reconstruct the context image in pixel space from frozen embeddings.
+The generation of diverse image variations via reconstruction can be attributed to two primary factors.
+  1. 1.
+    The frozen embeddings from the image encoder retain information only useful for the upstream task it is trained on.
+This in turn, leads to a lossy representation of the original image and the diffusion model has to fill in the missing details.
+  2. 2.
+    The pretrained diffusion decoder provides implicit regularization that prevents the model from simply collapsing to the conditional input image.
+We train a diffusion model from scratch to optimize Eq 2 with DINOv2 frozen embeddings.
+Qualitatively, we see that extremely early in training, less than 100K steps, the generated samples almost collapse to the conditional image.
+Fig.
+2 displays three samples from the trained diffusion model, showing some very minor low-level variations but no high-level variations.
+Similar results can be seen using SigLIP frozen embeddings in App.
+B.
+This suggests that pretraining or jointly training on a text-to-image pretraining objective may be the principal source of image variations.
+While this may be sufficient to generate reasonable image variations, it is non-trivial to predict the relationship between $f$ and the quality of image variations.
+For example, does a bigger text-to-image model lead to better image variations?
+Figure 2: A conditional diffusion model reconstructs images from frozen DINOv2 embeddings.
+Left: Input Images.
+Right: Three samples from the trained diffusion model with guidance 0.0 exhibiting low-level variation but lacking high-level variation.
+5 Image Variations via Web-Scale Image Pairs
+Inspired by web-scale image-text pretraining (Radford et al., 2021), we modify the denoising objective as follows:
+$\mathbb{E}_{t\sim\mathcal{U}(0,1),{\bm{{\epsilon}}}\sim\mathcal{N}(0,{\mathbf{I}})}\big{[}||{\bm{x}}-f({\bm{z}}_{t},t,{\bm{y}}_{c})||^{2}\big{]}$ (3)
+where ${\bm{x}}$ is an image from a webpage and ${\bm{y}}_{c}$ are frozen embeddings obtained from another random image from the same webpage.
+In particular, we use Episodic WebLI (Chen et al., 2023), where each episode contains randomly sampled loosely related images (i.e., they are clustered according to their URL).
+Note that Episodic Webli is explicitly deduplicated from all standard image train and test benchmarks.
+While Episodic Webli was originally designed for training few-shot vision language models, we introduce a novel application by utilizing it to train image variation models.
+We randomly sample an image as the conditioning input ${\bm{x}}$ and another image from the same episode as the ground-truth ”target” image ${\bm{y}}_{c}$.
+Each episode consists of images that are loosely related, whereas our model assumes conditioning and target images share semantic information.
+This mismatch may lead the model to waste capacity on modeling irrelevant noisy conditioning-target pairs.
+To address this we filter out pairs with low similarity as done in image-text pretraining.
+The pretrained encoder computes the global CLS representation from the conditioning and target image.
+We compute the cosine similarity between the global conditioning/target representations and filter out pairs below a lower threshold.
+Unlike image-text pretraining, we also filter conditioning-target pairs with similarity above a high threshold.
+in order to ensure generation of interesting images.
+This can ensure that generated images retain the semantics of the conditioning image, while being sufficiently different and interesting.
+App.
+D provide more information on how the high and low thresholds were set.
+After filtering, we obtain a dataset of around 50M image pairs.
+Notably, despite utilizing a dataset an order of magnitude smaller than standard text-to-image datasets (Schuhmann et al., 2021), we achieve strong results on image variations.
+6 Experiments
+6.1 Architecture Details
+For our baseline Semantica model, we inherit all hyper-parameters from the ImageNet label conditioned diffusion model.
+The denoising model follows a U-ViT architecture that operates on $256\times 256$ images.
+The architecture consists of a initial $1\times 1$ convolution.
+The model has four downsampling stages, where each stage downsamples the feature maps by a factor of 2 at its output and a final transformer stage.
+The resolution of the lowest feature map is $16\times 16$.
+Transformer blocks operate at the stages with the two lowest resolutions $16\times 16$ and $32\times 32$ and convolutional blocks operate in the remaining stages.
+The four downsampling stages have 128, 128, 256 and 512 channels and the final transformer has 1024 channels.
+The first three stages have three blocks each and the last stage has sixteen blocks.
+The optimizer is Adam (Kingma & Ba, 2014) with parameters $\beta_{1}=0.9$, $\beta_{2}=0.99$, $\epsilon=1e^{-12}$, batch size of 2048 and a learning rate of $2e^{-4}$.
+We also use Polyak averaging with a decay factor of $0.9999$.
+The diffusion loss parameters include v-prediction with loss in epsilon phase and a cosine adjusted schedule with a noise resolution of 32.
+We use the DDPM sampler with an interpolation of 0.2 (standard deviation is $\sigma_{t\to s}^{0.2}\sigma_{st}^{0.8}$) and $0.5$ guidance for our ablations.
+Each ablation run utilizes 256 TPUv3 (Google, 2023) chips around 300K steps.
+However, the consistent ranking of different ablations throughout training can allow for a much shorter training schedule to identify the best model.
+We report the FID on 50000 ImageNet samples.
+Our final Semantica model that operates on $512\times 512$ has a $2\times 2$ patchification layer instead of $256\times 256$.
+We then use 128, 256, 1024, 2048 and 4096 channels with 2, 3, 3, 3 and 12 blocks each.
+6.2 Choice of Image Encoder
+We first compare two choices of conditioning the diffusion model on frozen image embeddings.
+Global feature conditioning with FiLM layers and local feature conditioning with cross-attention.
+Fig.
+7 shows that cross-attention with local features, consistently outperforms FiLM across both SigLIP and DINO encoders.
+The result highlights the importance of local features for image variations.
+Then, we investigate the impact of scaling pretrained encoders and diffusion models.
+We evaluate eight combinations of two encoders each with Base and Large scales and two scales of diffusion models (SiD B and SiD L).
+Table.
+1 reports the FID of each of these combinations at 300K steps.
+Scaling the encoder while keeping the diffusion model fixed, offers improvements ranging from -0.8 for (SiD-L + DINO) to -1.5 for (SiD-B + SigLIP).
+Scaling the diffusion model with a fixed encoder size consistently improves FID by around -4.0 for all encoders.
+Finally, jointly scaling the encoder and diffusion model together results in significant improvements: DINOv2 improves from 13.0 to 9.0 and SigLIP from 17.1 to 11.2.
+Thus, from here on we will use DINOv2 as the frozen image encoder.
+SiD B SiD L
+DINOv2 B 13.0 9.8
+DINOv2 L 11.7 9.0
+SiD B SiD L
+SigLIP B 17.1 12.9
+SigLIP L 15.6 11.2
+Table 1: We report the FID on ImageNet across two encoders (DINOv2 and SigLIP), two diffusion model sizes (SiD B and SiD L) and two encoder sizes (B and L) at 300K steps.
+DINOv2 encoder performs better than SigLIP across all setups.
+Joint scaling of both the diffusion model and the image encoder works best for both setups.
+7 Metrics for Image Variations
+Previous methods assess image variations using the following approach.
+For each reference image in a test set, the model generates a new image.
+To quantify the quality of image variations, these methods employ LPIPS (Zhang et al., 2018) for individual image comparisons and FID (Heusel et al., 2017) or Precision/Recall (Kynkäänniemi et al., 2019) to compare test to generated distributions.
+These metrics can be sufficient to compare variations of our model on our episodic dataset since we explicitly filter out near duplicates, and in theory, the model is unlikely to repeat the same image.
+However, one major drawback is its inability to measure how diverse the generates samples are with respect to the input image.
+For instance, a model that just copies the input image or produces very minor variations will have near perfect LPIPS and FID.
+Thus, these metrics are not ideal for image variation baselines that rely on reconstruction.
+To address this limitation, we propose a few-shot approach to assess image variations.
+Given a set of $N$ test images, we randomly select $K$ images and generate $N/K$ samples for each, resulting in a total of $N$ samples.
+We then evaluate FID, recall and precision between the $N$ test images and $N$ samples.
+Precision measures the quality of the generated samples while recall measures sample diversity.
+Here we demonstrate how few-shot recall effectively captures the diversity of the generated samples.
+As a reminder, recall is the proportion of test images that lie within the manifold of generated samples.
+For each generated sample $i$, let $d_{i}$ be the distance to its $K$ th nearest neighbour.
+A ground-truth image has recall 1.0, if it lies within $d_{i}$ to any generated sample $i$.
+In the extreme case, where a model copies the same image during generation, $d_{i}=0$ for all $i$.
+Thus a new test image will have $d_{i}>0.0$ for all $i$, which leads to recall 0.0.
+Conversely, a sample that steers far away from the conditioning image is less likely to be represented in the real data manifold and is hence likely to have low precision.
+Finally, we also employ few-shot FID that provides a single score combining both quality and diversity.
+8 Comparisons
+8.1 Baselines
+We compare Semantica to state-of-the-art image variation baselines: Versatile Diffusion (Xu et al., 2023b), Stable Diffusion v2 Image Variations and IP-Adapter (Ye et al., 2023) on generating image variations of size $512\times 512$.
+As seen in Sec.
+2.1, all baselines rely on image reconstruction to generate image variations.
+We sweep across a range of guidance factors for all baselines.
+See App.
+E) for detailed results of FID with respect to guidance.
+8.2 ImageNet Oneshot
+Model One-Shot FID
+SD Image Variations v2 34.5
+Versatile Diffusion 26.3
+IP-Adapter 20.2
+Semantica 18.5
+Figure 3: Comparison of Semantica against three state-of-the-art image variation baselines on one-shot ImageNet, using evaluation metrics: FID (Left Table) and Precision-Recall (Right Plot:) as evaluation metrics.
+Each point in Fig.
+3 Right represents a different guidance factor.
+Semantica outperforms image-variation baselines achieving lower FID and a better precision-recall tradeoff.
+Setup.
+We compare Semantica with Versatile Diffusion, IP-Adapter and SD-v2 Image Variations on one-shot ImageNet.
+Remember that ImageNet has a total of thousand classes.
+We sample ten images randomly per-class and create a ground truth set of 10000 images.
+Each baseline model receives one image per-class and generates ten samples per-image with different random seeds, leading to a total of 10000 samples.
+We then compute FID, precision and recall between 10000 ground truth images and 10000 samples.
+Results.
+Fig.
+3 left reports the one-shot FID of all four models.
+For each model, we tune the guidance factors.
+Table 6 provides detailed results on the relationship between guidance factors and FID for each model.
+Semantica improves over the previous image variation models achieving a FID of 17.9, 2.3 over the second-best model IP-Adapter.
+Fig.
+3 right displays the precision-recall tradeoff of all models across different guidance factors.
+Since the conditioning information is present to all models, note that all models have much higher precisions than recall.
+IP-Adapter and Versatile Diffusion have precision greater than 0.8.
+SD-V2 IV has higher recall but much lower precision.
+At lower precisions, Semantica achieves similar recall as compared to SD-V2 IV with much higher precision.
+At a precision of 0.8, Semantica achieves a high recall higher than the IP-Adpater baseline.
+Semantica achieves the best-tradeoff between precision and recall among all the baselines.
+8.3 Label Grouped Baseline
+ImageNet Bedroom Church SUN397
+Label grouped 4.8 46.2 27.1 29.7
+Semantica 18.4 6.2 17.3 6.7
+Guidance @@@@ 0.5
+Label grouped 5.1 34.2 20.4 22.4
+Semantica 6.2 2.4 4.0 2.5
+Table 2: Comparison between Semantica and a Label Grouped baseline (LG) trained on ImageNet.
+The conditioning and target pairs have the same label for LG.
+LG outperforms Semantica in-distribution and performs worse on out-of-distribution datasets.
+Here we compare Semantica to a baseline that has direct label supervision (for example, ImageNet) on lower resolution images $256\times 256$.
+Recall that the conditioning and target image belong to the same webpage.
+However, in the presence of label supervision (as in ImageNet), the target and conditioning image can just belong to the same class label.
+So as a supervised baseline, we group images on ImageNet as per their label and train Semantica on this dataset.
+Table 2 compares the FID of the Label Grouped baseline (LG) to Semantica.
+Since LG is trained on ImageNet, it significantly outperforms Semantica (FID 4.8 vs FID 18.4).
+However, this trend reverses on all other datasets, where Semantica outperforms LG.
+Both the supervised baseline and Semantica rely on the DINOv2 encoder which was trained on a wide variety of data sources.
+Therefore the encoder itself may provide useful representations on a number of datasets.
+But training LG just on ImageNet, might limit the diffusion model’s exposure from non ImageNet images, potentially explaining its significant performance drop on all other datasets.
+9 Conclusion and Limitations
+Our paper explores a new method for training image-conditioned diffusion models to generate image variations.
+Instead of the typical image reconstruction approach, we condition the model on one random image from a webpage and train it to denoise another random image from the same webpage.
+Through rigorous evaluations, DINOv2 as the image encoder produces better image variations than the popular CLIP model.
+Finally, we emphasize the difficulty in measuring image variations, and propose new metrics that are applicable in the one-shot setting.
+In this work, we focus on evaluating image variations on datasets consisting of mainly a single object.
+When multiple objects are present in an image, additional supervision in the form of bounding boxes or text can allow for fine-grained control of generations.
+Further, as in prior works, we focus on frozen image encoders to efficiently encode representations and filter data as opposed to training an image encoder end-to-end.
+Thus Semantica can inherit the biases of the frozen image encoder.
+We leave studying the tradeoffs between finetuning and using frozen representations to future work.
+Appendix A More Samples: ImageNet
+Semantica
+IP-Adapter
+Semantica
+IP-Adapter
+Semantica
+IP-Adapter
+Semantica
+IP-Adapter
+Semantica
+IP-Adapter
+Figure 4: We present additional samples and comparisons on ImageNet. Samples from Semantica reflect diversity while being congruent with the conditioning image.
+Appendix B CLIP: Reconstruct Images
+We train a diffusion model conditioned on SigLIP embeddings to reconstruct the original image.
+Fig.
+5 shows four samples images from the ImageNet validation set and the corresponding generations from the generative model.
+Figure 5: A conditional diffusion model reconstructs images from frozen SigLIP embeddings.
+As seen in the case with frozen DINOv2 embeddings in Fig.2, the generated samples exhibit very minor low-level variations.
+Appendix C Data Filtering
+Figure 6: FID of DINO-v2 B/14 + Cross Attention with and without data filtering
+We investigate the impact of semantic data filtering.
+We note that DINOv2 models are more aggressive than CLIP in assigning low similarities to image pairs.
+So we set the lower threshold of CLIP and DINOv2 models such that, the total number of examples are roughly the same.
+This leads to a lower threshold of 0.65 for CLIP and 0.9 for DINOv2.
+We set the higher threshold to 0.9 in order to filter out near-duplicates.
+Fig.
+7 middle shows the FID of the DINO-v2 B/14 + Cross Attention with and without data filtering.
+Similarity-based data filtering in DINO feature space positively impacts the generation quality and improves FID by greater than 10.
+In future work, we can explore tuning these thresholds for a desired quality-diversity tradeoff or even directly conditioning the diffusion model on the desired similarity with the conditioning image.
+Appendix D Film vs Cross-attention
+Here, we compare Film based conditioning to cross-attention based conditioning.
+Figure 7: We plot ImageNet FID as a function of number of training steps on Episodic WebLI. Left: DINO-v2 B/14 with Film and cross attention Right: SigLIP B/14 with film and cross attention.
+Appendix E ImageNet one-shot FID vs Guidance Hyperparameters
+We report fine-grained FID results for different guidance values. Tab. 3 reports fine-grained FID results for SD-v2 IV and Versatile Diffusion. Tab. 5, reports results for IP-Adapter and Semantica.
+Guidance FID
+1.0 46.7
+4.0 30.8
+6.0 34.5
+8.0 37.6
+Guidance FID
+1.0 28.5
+4.0 26.3
+6.0 29.4
+8.0 31.8
+Table 3: Guidance against one-shot ImageNet FID. Left: SD-IV and Right: Versatile Diffusion
+Guidance Scale FID
+1.0 0.5 57.4
+4.0 0.5 25.4
+7.0 0.5 24.5
+1.0 1.0 42.6
+2.0 1.0 23.2
+4.0 1.0 20.2
+7.0 1.0 20.6
+7.0 1.0 21.2
+Guidance FID
+0.1 29.4
+0.5 21.0
+1.0 18.5
+Table 4: Semantica
+Table 5: Guidance against one-shot ImageNet FID. Left: IP Adapter and Right: Semantica
+Appendix F SUN-397 Oneshot
+This experiment compares Semantica with IP-Adapter on one-shot SUN-397.
+SUN-397 has a total of 397 classes.
+We sample 25 images randomly per-class and create a ground truth set of 9925 images.
+Each model generates 25 samples given a randomly sampled image, leading to total of 9925 samples.
+Similar to ImageNet, Fig.
+8 reports FID, precision and recall between the 9925 generated samples and ground-truth images.
+Semantica achieves a one-shot FID of 13.0, outperforming IP-Adapter.
+It also achieves a much more favourable precision-recall tradeoff.
+Model One-Shot FID
+IP-Adapter 12.8
+Semantica 12.6
+Figure 8: Comparison of Semantica against IP-Adapter on one-shot SUN397, using evaluation metrics: FID (Left Table) and Precision-Recall (Right Plot:) as evaluation metrics.
+Semantica
+IP-Adapter
+Semantica
+IP-Adapter
+Figure 9: We present additional samples and comparisons on SUN397
+Guidance Scale FID
+1.0 0.5 24.6
+4.0 0.5 14.1
+7.0 0.5 16.7
+1.0 1.0 20.0
+4.0 1.0 17.8
+7.0 1.0 27.9
+Guidance FID
+0.1 13.2
+0.3 12.6
+0.5 13.0
+0.7 13.5
+1.0 14.7
+Table 6: Guidance against one-shot SUN397 FID. Left: IP Adapter and Right: Semantica
+Appendix G Qualitative Effect of Guidance
+Fig.
+10, displays five conditioning images from ImageNet and the generated samples at different guidance factors.
+At guidance factor 0.0, the samples reflect a broad semantic category from the conditioning image.
+Increasing the guidance factor leads to samples that incorporate more specific details from the conditioning image.
+For example, with the conditioning image of the dog and the kid, Semantica stars with a sample of a dog.
+The specific breed of the dog and the child in the image appear as we amplify the guidance.
+Fig.
+11 showcases samples for each small dataset across various guidance factors.
+In row four, the bed (main object) persists across all guidance levels, while the chair and the fence appear at high guidance levels.
+Row five exhibits a similar effect: the number of minarets in the generated church increases from one to two and the shape of the main dome begins to resemble the conditioning image.
+In row one, the sample resembles toys with zero guidance, the sample resembles toys but transforms into a crowded convention as guidance increases.
+Figure 10: Left: Conditioning Image from ImageNet.
+Right: Generated samples with guidance factors 0.0, 0.1, 0.2, 0.5 and 1.0.
+At guidance factor 0.0, the samples reflect a broad semantic category from the conditioning image.
+Increasing the guidance factor leads to samples that incorporate more specific details from the conditioning image.
+SUN397
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+LSUN Bedrooms
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+LSUN Churches
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+Figure 11: Left: Conditioning Images from SUN397 (Top two rows), LSUN Bedrooms (Middle two rows) and LSUN churches (Last two rows).
+Right: Generated samples with guidance factors 0.0, 0.1, 0.2, 0.5, 1.0 and 1.5.
 ImageNet
@@
 Semantica requires no label supervision or finetuning.
-Machine learning initially focused on optimizing and improving models on small datasets.
-The field has transitioned to training general purpose models on web-scale data and then finetuning them for specific tasks on smaller datasets.
-This paradigm shift has lead to state-of-the-art results on a number of different domains.
-In this paper, we focus on the relatively underexplored task of adapting an image generative model to different datasets.
-One approach is to simply train a generative model on a large dataset of unlabelled images and finetune them on smaller datasets.
-While this approach is straight-forward in theory, it requires clever architecture or regularizer design to prevent overfitting in practice (See Sec.2.1).
-As models scale up, finetuning for every dataset also just becomes increasingly impractical.
-Instead, we apply the principle of in-context learning to adapt a image generative model.
-We train our model called Semantica using image episodes, which are image pairs that belong to the same web page.
-Therefore, training relies exclusively on the hypothesis that images from the same web page should have some common semantic attributes.
-For example, it is highly likely that images scraped from a Wikipedia page on dolphins, contain pictures of dolphins.
-At adaptation, the model receives an image and then generates another image that preserves semantic information.
-Semantica consists of a pre-trained image encoder and a diffusion model.
-The encoder extracts representations from the conditioning image and guides the diffusion model.
-We thoroughly investigate the impact of different choices for the image encoder (contrastive or self-supervised) and the conditioning mechanisms (global or local).
-Our experiments also highlight the importance of semantic data filtering to generate high-quality images.
-Each row in Fig 1 showcases a conditioning image to the left and five generated samples to the right.
-Semantica is capable of generating high quality and diverse images, reflective of semantic information from the conditioning image across a number of different datasets.
-2 Related work
-2.1 Generative Transfer
-There has been extensive research that studies the adaptation of source-pretrained generative models to a target dataset with adaptation of weights.
-Initial works study the transfer of discriminators and generators in GANs from a source dataset to a target dataset [47].
-Further, [17] show that ImageNet pretraining on a large GAN model is beneficial for transfer to small datasets.
-A number of papers focus on improving generation quality by adapting only a subset of parameters.
-These include scale and shift parameters [32], updating only the higher discriminator layers [28], linear combinations of scale and shift parameters [41], modulating kernels or convolutions [59, 58, 10, 2] and singular values [38], mapping networks from noise to latents [46, 29, 53] and latent offsets [12].
-Various works apply regularization losses by enforcing constraints to samples/weights by the source generator including elastic weight regularization [27], domain correspondence [33, 16, 22], contrastive learning [60], spatial alignment [51], inversion [49, 23, 44], random masks on discriminators [61] and alignment free spatial correlation [30].
-Given the increasing popularity of VQ-VAE and diffusion based models, recent work [43] and [61] explore few-shot finetuning on VQ-VAE tokens and diffusion models.
-We defer to the survey [1] for a detailed exposition of all these methods.
-In contrast to these works, we explore training a generator on web-scale images and study their transfer to standard small-scale image datasets.
-Retrieval augmented models [7, 5] compute nearest neighbours for a query image across a bank of memory images.
-These retrieved neighbors facilitate the training or generation process.
-Unlike these methods, we do not require access to a memory bank of images during train or test time.
-2.2 Diffusion
-Diffusion and score-based generative models have become increasingly successful in modelling images [20, 40, 31, 4], videos [18, 42] and audio [25].
-As generation quality has steadily improved, they have been used in contexts with more and more conditioning variables.
-Well-known examples are text-to-image and text-to-video modelling, where the conditioning variable is text.
-In this case, the conditioning variable can be seen as a sequence from which cross-attention layers communicate to the feature maps of the image or video, i.e.
-what the diffusion model is learning to generate [40, 31].
-As the desire for controllable generation increases, solutions such as ControlNet [56] have been developed.
-ControlNet takes in conditioning images of the same size as the generations, and uses a copy of the UNet to learn an encoder for the conditioning signals.
-Although this encoder trains fast due to parameter initialization from a pretrained diffusion UNet, it is difficult to deal with different sized inputs.
-In those cases, only conditioning via cross-attention as done in [52] overcomes the in-place additions between the ControlNet encoder and the base UNet.
-Conditioning on images as context has produced impressive results, turning scribbles or edge detections into high quality image generations [48] and discriminative tasks [3, 26].
-In contrast with the above mentioned techniques, our framework relies on general web-based pretraining for semantic-based adaptive image generation.
-Finally, while [13] study the transfer of few-shot diffusion models between small datasets (CIFAR-100 -> miniImageNet, we see in Sec.
-5.6 that this can lead to sub-optimal transfer.
-3 Background
-3.1 Diffusion
-Diffusion models learn to generate examples by gradually denoising a diffusion process.
-For a single datapoint, their loss can be expressed as a squared error between the original datapoint and its prediction:
-$\mathbb{E}_{t\sim\mathcal{U}(0,1),{\bm{{\epsilon}}}\sim\mathcal{N}(0,{\mathbf{I}})}\big{[}w(t)||{\bm{x}}-f({\bm{z}}_{t})||^{2}\big{]}\text{ where }{\bm{z}}_{t}=\alpha_{t}{\bm{x}}+\sigma_{t}{\bm{{\epsilon}}}_{t}$ (1)
-It is helpful to define $\mathrm{SNR}(t)=\alpha_{t}^{2}/\sigma_{t}^{2}$.
-In the case of $w(t)=\mathrm{SNR}(t)$ the loss above is equivalent to a loss in ${\bm{{\epsilon}}}$-space, the simple loss from [19].
-After training, the denoising model generates samples by taking small steps.
-Starting at $t=1$ with initial Gaussian noise and one slowly denoises for timesteps $t=1,1-1/N,\ldots$ where the number of sampling steps is $N$.
-Although many samplers are possible, in this paper we use the standard DDPM sampler [19].
-Architecture
-In early days, diffusion literature typically used UNets that consisted of ResNet blocks, with optional self-attention layers.
-More recent architecture either use full Transformers (DiT [35], StableDiffusion) or transformer backbones UViTs [21].
-This transformer backbone makes it especially easy to use conditioning signals in these architecture via cross-attention layers.
-To be precise, one can say that the denoising neural network takes in ${\bm{z}}_{t}\in\mathbb{R}^{H\times W\times 3},t\in\mathbb{R},y\in\mathbb{R}^{T_{y}\times D_{y}}$.
-In principle it does not matter which diffusion model we use to generate images.
-In fact, every generative model would suffice.
-In practice we use the simple diffusion framework [21] because it can learn to generate high resolution images end-to-end without the need of a separate autoencoder.
-4 Method
-4.1 Episodic Dataset
-The training dataset for our model is a subset of WebLI [9] that consists of roughly one billion images from the web.
-In particular, we use Episodic WebLI [8], where each episode contains randomly sampled loosely related images (i.e., they are clustered according to their URL).
-We randomly sample an image as the conditioning input and another image from the same episode as the ground-truth "target" image.
-We do not use any text inputs and only rely on image conditioning.
-4.2 Image Encoder
-Training an image-conditioned diffusion model requires two components: 1) An image encoder that extracts semantic information from a conditioning image.
-2) A diffusion model that denoises a noisy target image conditioned on the image representations.
-We could train a separate Vision Transformer or ResNet model end-to-end as an image encoder with the diffusion model to learn useful conditioning representations.
-Instead, we leverage pre-trained image encoders and condition our diffusion model on their “frozen" representations.
-This offers two advantages.
-First, we can precompute representations for all images in the dataset that eliminates expensive forward passes through the encoder during training.
-Second, by relying on pretrained encoders, we can just focus on scaling the diffusion model.
-We investigate ViT image encoders trained with two pretraining strategies: training on web-text image pairs [37] (we use the SigLIP [55]), and self-supervision with teacher pseudo-labels (DINO) [6, 34].
-Furthermore, we explore two conditioning strategies to incorporate conditioning information:
-  1. 1.
-    Token level conditioning: The diffusion model cross-attends to a sequence of tokens from the final transformer block of the image encoder.
-  2. 2.
-    Global feature conditioning: The encoder provides a single global feature representation from the encoder, being the CLS token for DINO and the output of the MAP block for SigLIP.
-We normalize and embed this with a dense projection and add it to the noise embedding.
-This subsequently conditions the diffusion model via FiLM layers [36] as done in standard label-conditioned diffusion.
-4.3 Data Filtering
-Each episode consists of images that are loosely related, whereas our model assumes conditioning and target images share semantic information.
-This mismatch may lead the model to waste capacity on modeling irrelevant noisy conditioning-target pairs.
-To address this, we filter out pairs with low semantic similarity.
-The pretrained encoder (DINO/SigLIP) computes the global CLS representation from the conditioning and target image.
-We compute the cosine similarity between the global conditioning/target representations and filter out pairs below a lower threshold.
-Additionally, to ensure generation of interesting images, we also filter conditioning-target pairs with similarity above a high threshold.
-This may ensure that generated images retain the semantics of the conditioning image, while being sufficiently different.
-5 Experiments
-5.1 Evaluation Datasets
-We use $10\%$ of the Episodic WebLI dataset for the architecture design ablations and we use the entire dataset for the final model.
-We evaluate the model by conditioning it on ImageNet validation set images and then computing the FID between the generated samples and the ImageNet validation set.
-We investigate the transfer of Semantica on four datasets: ImageNet [39], LSUN-Outdoor Church, LSUN-Bedroom [54] and SUN-397 [50].
-We report the FID for each dataset using 50K conditioning images randomly sampled from within the dataset.
-We use the validation dataset for ImageNet while the train set for all other datasets.
-To assess diversity, we compute the mean LPIPS [57] score between 50K generated samples and their corresponding conditioning images.
-5.2 Image Encoder
-Here, we evaluate the different encoder and conditioning design choices (Sec.
-4.2).
-We train baseline diffusion models with two frozen image encoders (DINO-v2 B/14 and SigLIP B/16) and two conditioning strategies (cross-attention and FILM).
-Refer to captionRefer to captionRefer to caption
-Figure 2: We plot ImageNet FID as a function of number of training steps on Episodic WebLI.
-Left: FID of two image encoders DINO-v2 B/14 and SigLIP B/16 using two different conditioning methods: 1) All encoder output tokens via cross-attention (CA) 2) Global feature representation via FILM.
-Center: FID of DINO-v2 B/14 + Cross Attention with and without data filtering.
-Right: FID of DINO-v2 B/14 + Cross Attention with 1x1 patching and higher learning rate.
-Fig.
-2 left shows ImageNet FID as a function of number of training steps on Episodic WebLI.
-Token-level conditioning significantly outperforms global feature conditioning for both DINO-v2 and SigLIP, highlighting the importance of low-level features for this particular task.
-While SigLIP FiLM performs better than DINO-v2 FiLM, this trend reverses with cross-attention.
-This suggests a greater compatibility of DINO-v2 low-level features over CLIP low-level features for semantics-based image generation.
-For this reason, we will now use DINO-v2 frozen representations with a cross-attention conditioning mechanism as our main architecture.
-5.3 Data Filtering
-We investigate the impact of semantic data filtering, as described in Sec.
-4.3.
-We heuristically set the lower and higher thresholds to 0.3 and 0.9 respectively.
-Fig.
-2 middle shows the FID of the DINO-v2 B/14 + Cross Attention with and without data filtering.
-Similarity-based data filtering in DINO feature space positively impacts the generation quality and improves FID by greater than 10.
-In future work, we can explore tuning these thresholds for a desired quality-diversity tradeoff or even directly conditioning the diffusion model on the desired similarity with the conditioning image.
-5.4 Hyperparameters
-For our baseline diffusion model, we inherit all hyper-parameters from the ImageNet label conditioned diffusion model.
-The denoising model follows a U-ViT architecture that operates on $256\times 256$ images.
-The architecture consists of a initial $2\times 2$ patchification layer [11, 45] that downsamples the noised image by a factor of 2.
-The model has four downsampling stages, where each stage downsamples the feature maps by a factor of 2 at its output.
-The resolution of the lowest feature map is $16\times 16$.
-Transformer blocks operate at the stages with the two lowest resolutions $16\times 16$ and $32\times 32$ and convolutional blocks operate in the remaining stages.
-The four downsampling stages have 128, 256, 512 and 1024 channels.
-The first three stages have three blocks each and the last stage has sixteen blocks.
-The optimizer is Adam [24] with parameters $\beta_{1}=0.9$, $\beta_{2}=0.99$, $\epsilon=1e^{-12}$, batch size of 2048 and a learning rate of $1e^{-4}$.
-We also use Polyak averaging with a decay factor of $0.9999$.
-The diffusion loss parameters include v-prediction with loss in epsilon phase and a cosine adjusted schedule with a noise resolution of 32.
-We use the DDPM sampler with an interpolation of 0.2 (standard deviation is $\sigma_{t\to s}^{0.2}\sigma_{st}^{0.8}$) and no guidance for our ablations.
-Each ablation run utilizes 256 TPUv3 [15] chips around 2 days.
-However, the consistent ranking of different ablations throughout training can allow for a much shorter training schedule to identify the best model.
-We finally perform a lightweight sweep of the diffusion model hyper-parameters.
-Notably, two minor modifications improve generation quality: 1) Replacing the 2x2 patching with no patching.
-The model consists of a $1\times 1$ convolution at the input and output.
-The downsampling proceeds with an additional stage consisting of two convolutional blocks and 128 channels each.
-2) Increasing the learning rate from $41e-$ 4 to $42e-$ 4.
-Fig.
-2 right shows the improvement in FID of the DINO B/14 encoder with these modifications.
-Other changes such as tuning noise schedules, dropout rates or Polyak decay factors cause negligible changes or decrease sample quality.
-5.5 Final Training
-In our final training phase, we make the following four adjustments:
-  1. 1.
-    Replace DINO-v2 ViT-B/14 conditioning features with the larger DINO ViT-L/14 conditioning features.
-  2. 2.
-    Scale the model by doubling the number of channels across all UViT stages. The only exception is the compute-heavy first downsampling stage which we keep at 128.
-  3. 3.
-    Expand training data to include the entirety of episodic WebLI which was previously $10\%$ in the ablations.
-  4. 4.
-    Extend the training duration from 300K steps to two million steps
-We train our model on 256 Cloud TPUv5 chips [14] for approximately 12 days with a constant learning rate of $2\text{e}{-4}$.
-5.6 Label Grouped Baseline
-Here we compare Semantica to a baseline that has direct label supervision (for example, ImageNet).
-Recall that the conditioning and target image belong to the same webpage.
-However, in the presence of label supervision (as in ImageNet), the target and conditioning image can just belong to the same class label.
-So as a supervised baseline, we group images on ImageNet as per their label and train Semantica on this dataset.
-Table 1 compares the FID of the Label Grouped baseline (LG) to Semantica.
-Since LG is trained on ImageNet, it significantly outperforms Semantica (FID 4.8 vs FID 18.4).
-However, this trend reverses on all other datasets, where Semantica outperforms LG.
-Both the supervised baseline and Semantica rely on the DINOv2 encoder which was trained on a wide variety of data sources.
-Therefore the encoder itself may provide useful representations on a number of datasets.
-But training LG just on ImageNet, might limit the diffusion model’s exposure from non ImageNet images, potentially explaining its significant performance drop on all other datasets.
-5.7 One Shot Results
-We evaluate Semantica’s one-shot image generation capabilities.
-Each row in Fig.
-1 showcases a conditioning image to the left and five generated samples to the right.
-We observe that Semantica is capable of generating high quality and diverse images.
-Table 1 reports quantitative results.
-Semantica receives 1000 and 397 conditioning images for ImageNet and SUN397 respectively, one per class.
-The model generates 50 and 120 samples per class for ImageNet and SUN397 totalling 50000 and 47840 generated samples.
-We then report the FID with 50000 ground truth samples.
-Semantica achieves a FID of 19.1 and 9.5 on ImageNet and SUN397 respectively.
-On increasing the guidance to 0.5, the FID improves to 9.3 and 7.1 respectively.
-Simply replicating the conditioning image Copy 50 and 120 times, achieves a much worse FID of 38.4 and 70.6.
-5.8 Effect of Guidance
-We investigate the quality/diversity tradeoff offered by Semantica’s guidance parameter.
-Fig.
-3 Left illustrates the FID and LPIPS on ImageNet for different guidance factors.
-Note that we filter out near duplicates with a semantic similarity greater than 0.9 from the training set.
-This makes it unlikely that the model copies the conditioning image exactly to achieve a perfect FID.
-With no guidance, Semantica achieves a FID of 18.4 on ImageNet.
-The high LPIPS score 0.72 further reflects the model’s ability to generate diverse samples.
-On increasing the guidance upto a certain extent, we tradeoff quality with diversity.
-Notably, at guidance factor 0.5, which is commonly used in label conditioned generation, Semantica achieves a FID of 6.2 with a LPIPS score of 0.68.
-Fig.
-4, displays five conditioning images from ImageNet and the generated samples at different guidance factors.
-At guidance factor 0.0, the samples reflect a broad semantic category from the conditioning image.
-Increasing the guidance factor leads to samples that incorporate more specific details from the conditioning image.
-For example, with the conditioning image of the dog and the kid, Semantica stars with a sample of a dog.
-The specific breed of the dog and the child in the image appear as we amplify the guidance.
-Refer to captionRefer to caption
-Figure 3: LPIPS (Higher is more diverse) vs FID (Lower is better) for five different datasets (Left: ImageNet, Right: LSUN Bedroom, LSUN Churches, SUN397).
-Each point represents a different guidance factor.
-As expected, we tradeoff quality with diversity on increasing guidance, but even with lower diversity the model does not collapse completely to its conditioning image.
-ImageNet Bedroom Church SUN397
-Label grouped 4.8 46.2 27.1 29.7
-Semantica 18.4 6.2 17.3 6.7
-Guidance @@@@ 0.5
-Label grouped 5.1 34.2 20.4 22.4
-Semantica 6.2 2.4 4.0 2.5
-ImageNet SUN397
-Semantica 19.1 9.5
-+ Guidance @@@@ 0.5 9.3 7.1
-Copy 38.4 70.6
-Table 1: Left: Comparison between Semantica and a Label Grouped baseline (LG) trained on ImageNet.
-The conditioning and target pairs have the same label for LG.
-LG outperforms Semantica in-distribution and performs worse on out-of-distribution datasets.
-Right (One Shot Generation Results): Evaluates Semantica one-shot image generation capabilities.
-Semantica receives 1000 and 397 conditioning images for ImageNet and SUN397 respectively, one-per-class.
-The model generates 50 and 120 samples per-class for ImageNet and SUN397 totalling 50000 and 47840 generated samples.
-We then report the FID with 50000 ground truth samples.
-Copy is a baseline where the same per-class conditioning image is just replicated 50 and 120 times respectively.
-Figure 4: Left: Conditioning Image from ImageNet.
-Right: Generated samples with guidance factors 0.0, 0.1, 0.2, 0.5 and 1.0.
-At guidance factor 0.0, the samples reflect a broad semantic category from the conditioning image.
-Increasing the guidance factor leads to samples that incorporate more specific details from the conditioning image.
-Fig.
-3 Right demonstrates quantitative results on three smaller datasets LSUN Bedroom, LSUN Church and SUN-397.
-We observe a similar tradeoff between quality and diversity as we vary the guidance factor.
-With zero guidance, Semantica achieves a FID of 6.2, 17.3 and 6.7 on LSUN Bedroom, LSUN Church and SUN397 respectively.
-On increasing the guidance to 0.5, the FID improves to 2.4, 4.0 and 2.5 respectively.
-Fig.
-5 showcases samples for each small dataset across various guidance factors.
-In row four, the bed (main object) persists across all guidance levels, while the chair and the fence appear at high guidance levels.
-Row five exhibits a similar effect: the number of minarets in the generated church increases from one to two and the shape of the main dome begins to resemble the conditioning image.
-In row one, the sample resembles toys with zero guidance, the sample resembles toys but transforms into a crowded convention as guidance increases.
-SUN397
-Refer to caption Refer to caption
-Refer to caption Refer to caption
-LSUN Bedrooms
-Refer to caption Refer to caption
-Refer to caption Refer to caption
-LSUN Churches
-Refer to caption Refer to caption
-Refer to caption Refer to caption
-Figure 5: Left: Conditioning Images from SUN397 (Top two rows), LSUN Bedrooms (Middle two rows) and LSUN churches (Last two rows).
-Right: Generated samples with guidance factors 0.0, 0.1, 0.2, 0.5, 1.0 and 1.5.
-6 Conclusion
-We present Semantica, an image-conditioned diffusion model capable of generating high-quality and diverse images based on the semantics of a conditioning image.
-Our pretraining strategy is simple, the diffusion model models a target image with the conditioning image coming from the same webpage.
-Our experiments also further emphasize the expressivity of pretrained image encoders and necessity of semantic-based data filtering.
-Limitations
-We identify the following limitations of Semantica.
-1) While adapting Semantica to downstream datasets, requires no finetuneing, a limitation is that training Semantica requires high computational resources as is the case with most generative models.
-However, any improvements that make diffusion models more efficient is likely to transfer to Semantica.
-2) Our model depends on a frozen encoder and we do not modify the weights.
-With more resources, we can also study the tradeoffs with this design decision.
-3) With a very high guidance factor (>1.0) , Semantica might converge to an oversaturated conditioning image.
-An area of improvement would be to find ways to steer it to another image with the same semantic information.
-4) Our model as is does not incorporate additional conditioning signals.
-Generalizing the model to utilize additional conditioning information can be an exciting area of future research.
-5) While our model generally produces diverse and high-quality images, sometimes artifacts can appear especially at guidance 0.0 and in areas like lower-level structure and human faces.
-Scaling our model further may remove these artifacts.
-Appendix A Broader Impact Statement
-This paper introduces a model capable of generating images that reflect the semantic information of a given input image.
-Potential positive applications include being a creative aid, augmentation of an image retrieval system or enhancing the size of an image dataset.
-While the model lacks direct controllability thus mitigating some negative applications like creation of personalized deepfakes, our model can still generate deepfakes.
-Therefore, we advocate for responsible usage for our model.
-Appendix B Standard Errors
-In general, the variance in FID is small.
-We compute the variance across different random seeds and different randomly sampled one shot datasets on ImageNet.
-With a fixed one-shot dataset and five different random seeds, we report a FID of $9.4\pm 0.02$.
-We also sample five different one-shot datasets and report a FID of $9.5\pm 0.06$.
-Appendix C ImageNet Samples
-Figure 6: Set 1: Conditioning Images from ImageNet
-Figure 7: Set 1: Random Samples 1
-Figure 8: Set 1: Random Samples 2
-Figure 9: Set 1: Random Samples 3
-Figure 10: Set 2: Conditioning Images from ImageNet
-Figure 11: Set 2: Random Samples 1
-Figure 12: Set 2: Random Samples 2
-Figure 13: Set 2: Random Samples 3
-Figure 14: Set 3: Conditioning Images from ImageNet
-Figure 15: Set 3: Random Samples 1
-Figure 16: Set 3: Random Samples 2
-Figure 17: Set 3: Random Samples 3
-Appendix D LSUN Bedroom Samples
-Figure 18: Conditioning Images from LSUN Bedroom
-Figure 19: Random Samples 1
-Figure 20: Random Samples 2
-Figure 21: Random Samples 3
-Appendix E SUN 397 Samples
-Figure 22: Conditioning Images from SUN397
-Figure 23: Random Samples 1
-Figure 24: Random Samples 2
-Figure 25: Random Samples 3
-Appendix F LSUN Church Samples
-Figure 26: Set 1: Conditioning Images from LSUN Churches
-Figure 27: Random Samples 1
-Figure 28: Random Samples 2
-Figure 29: Random Samples 3